OpenStudy (anonymous):

solve for n -4n+5/7= -1

OpenStudy (anonymous):

n=4/7

OpenStudy (anonymous):

n=4/7

OpenStudy (anonymous):

n-4n+5/7=-1 n-4n=-1-5/7 n(1-4)=-12/7 n=(-12/7)/(-3) n=4/7